OpenStudy (anonymous):

What is the equation of the hyperbola with foci at (0, ±10) and vertices at (0, ±8)?

OpenStudy (abb0t):

center(0,0)vertical transverse axis a=8 c=10 \[c^2-a^2=b^2\] b= ___ (y)² /36 - (x)²/ 64 =1 \[\frac{ y^2 }{ b^2 }-\frac{ x^2 }{ c^2 }=1\] (vertices (0,±a)